kuz:

Solve the system of equations: 2x + y = 10 x - y = 2

Anapooh:

mad easy I think YK how to do ts gng

kuz:

nah I actually dont

Anapooh:

you got ts from Google anyway

tax:

First, you'll add the equations: (2x + y) + (x - y) = 10 + 2 3x = 12 , then you'll divide which makes it x = 4 Finally, you'll substitute 4 into the equation using x - y = 2: 4 - y = 2 y = 2 So your final solution will be (x,y) = (4,2)

kuz:

thank you

Anapooh:

..

kuz:

bro ana syfm